OpenStudy (anonymous):

Study the following reaction carefully. What classification should this reaction have? 2Al + 3NiBr2 2AlBr3 + 3Ni

OpenStudy (kc_kennylau):

Single replacement reaction

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

A reaction is monitored and the products are found to be a salt and water. The reactant(s) must have been? a metal-nonmetal compound an acid and a base a metal and a metal-nonmetal compound two metal-nonmetal compounds

OpenStudy (kc_kennylau):

acid and base

OpenStudy (anonymous):

If approximately equal volumes of an acid and a base of equal strengths are combined, the resulting solution should have a pH of around 7 above 7 below 7 impossible to determine

OpenStudy (kc_kennylau):

\[\approx7\]

OpenStudy (anonymous):

≈7 whats that

OpenStudy (kc_kennylau):

around 7 sorry for late reply was browsing

OpenStudy (anonymous):

In the following reaction, which substance is the precipitate? (NH4)2SO4(aq) + Ba(NO3)2(aq) BaSO4(s) + 2NH4NO3(aq)

OpenStudy (kc_kennylau):

sorry I was browsing

OpenStudy (kc_kennylau):

Barium sulphate (it's the only not aq)

OpenStudy (kc_kennylau):

@kdday here? :)

OpenStudy (anonymous):

so its the Ba9no302(aq)

OpenStudy (kc_kennylau):

\(BaSO_4\)

OpenStudy (anonymous):

Study the following reaction carefully. What classification should this reaction have? H2SO4 + 2NaOH Na2SO4 + 2H2O synthesis decomposition single replacement neutralization

OpenStudy (kc_kennylau):

neutralization acid+base is always neutralization

OpenStudy (kc_kennylau):

word equation: Sulfuric acid+Sodium hydroxide->Sodium sulfate+Water

OpenStudy (anonymous):

okay i get it now, so When a double displacement reaction takes place, one of the products must be? a salt a gas a precipitate water

OpenStudy (kc_kennylau):

salt, because it's ionic :)

OpenStudy (anonymous):

i got a 99 yay! are you good at math too? i needa tutor like you

OpenStudy (kc_kennylau):

yep I am now helping others at math that's why i always have late reply :)

OpenStudy (kc_kennylau):

send the link to me after you ask the question :)

OpenStudy (abb0t):

He's not really tutoring, he's giving you answers and you're not learning much.

OpenStudy (anonymous):

okay abb0t and ur point is ! after i get the awnser i research and reaad !

OpenStudy (kc_kennylau):

i'm not just giving answers :)

OpenStudy (abb0t):

3.0 student all thanks to kind users like the one above giving you answers to your exams and quizzes.

OpenStudy (kc_kennylau):

thx :)

OpenStudy (abb0t):

I don't see "UE" anywhere on the top 100: http://www.forbes.com/top-colleges/

OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (2, 2) and perpendicular to the line whose equation is y = x y = x - 4 y = x + 4 y = -x + 4

OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (2, 2) and perpendicular to the line whose equation is y = x y = x - 4 y = x + 4 y = -x + 4

OpenStudy (kc_kennylau):

So they're perpendicular, which means that the product of the slopes is -1.

OpenStudy (kc_kennylau):

and the slope of the new line is therefore -1.

OpenStudy (kc_kennylau):

@kdday

OpenStudy (kc_kennylau):

substitute the coordinates into the equation :)

OpenStudy (anonymous):

@kc_kennylau okay so is it y = x - 4 y = x + 4 y = -x + 4